OpenStudy (anonymous):

Which one of these equations is correctly written in Standard Form? 1/3x + y = −4 3x + 2.5y = −4 x + 3y = −12 −3x + y = −12

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (akashdeepdeb):

The Standard Form equation of a line has the following formula: Ax + By = C where A ≠0 and B ≠0

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

are you sure???

OpenStudy (anonymous):

but its a multipul choice problem maybe i typed something wrong

OpenStudy (akashdeepdeb):

Yes check it again.

OpenStudy (anonymous):

Which one of these equations is correctly written in Standard Form? 1/3 x + y = −4 3x + 2.5y = −4 x + 3y = −12 −3x + y = −12

OpenStudy (anonymous):

no i copy and pasted it this time

OpenStudy (anonymous):

the third one

OpenStudy (anonymous):

tank you kelliegirl33

OpenStudy (anonymous):

you do not want x negative, you do not want decimals or fractions.

OpenStudy (anonymous):

YW :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

read this... http://www.algebralab.org/studyaids/studyaid.aspx?file=algebra1_5-5.xml

OpenStudy (akashdeepdeb):

Thanks! I think my mistake was I was overlooking the Fractions......